OpenStudy (anonymous):

24 divided 2 to the second power x(x-4)

OpenStudy (anonymous):

\[24\div2^{2}x(x-4)

OpenStudy (anonymous):

Your code needs to be rewritten please. I think it is just missing a closing

OpenStudy (anonymous):

\[24\div2^{2}x(x-4)\]

OpenStudy (anonymous):

\[\frac{ 24 }{ 4x(x-4) }\]?\[\frac{ 6 }{ x(x-4) }\]

OpenStudy (anonymous):

distribute the x x(x-a) = x*x - x*a

OpenStudy (anonymous):